A first order reaction has a half life period of 69.3s at 0.01mol literinverse reactant concentration ,rate wii be

Grade:9

1 Answers

Sai bharath
6 years ago
T1/2 = 69.3ST1/2=0.693/k (formula )K = 0.693/69.3 SK=0.01S^-1For a first order reaction Rate = k× concentration of reactantRate = 0.01 ×0.01mol L^-1 = 10 ^-4 M S^-1